# Mix Q-learning for Lane Changing: A Collaborative Decision-Making Method in Multi-Agent Deep Reinforcement Learning

### v1

@@ Abstract
Lane-changing decisions, which are crucial for autonomous vehicle path planning, face practical challenges due to rule-based constraints and limited data. Deep reinforcement learning has become a major research focus due to its advantages in data acquisition and interpretability. However, current models often overlook collaboration, which affects not only impacts overall traffic efficiency but also hinders the vehicle’s own normal driving in the long run. To address the aforementioned issue, this paper proposes a method named Mix Q-learning for Lane Changing(MQLC) that integrates a hybrid value Q network, taking into account both collective and individual benefits for the greater good. At the collective level, our method coordinates the individual Q and global Q networks by utilizing global information. This enables agents to effectively balance their individual interests with the collective benefit. At the individual level, we integrated a deep learning-based intent recognition module into our observation and enhanced the decision network. These changes provide agents with richer decision information and more accurate feature extraction for improved lane-changing decisions. This strategy enables the multi-agent system to learn and formulate optimal decision-making strategies effectively. Our MQLC model, through extensive experimental results, impressively outperforms other state-of-the-art multi-agent decision-making methods, achieving significantly safer and faster lane-changing decisions.
Index Terms:
Lane change , multi-agent deep reinforcement learning , automated vehicle
@@ I Introduction
In the field of CVIS, autonomous driving and related applications have the potential to revolutionise transport which could fundamentally change traffic patterns and provide significant convenience[2, 3, 4, 5]. Among the tasks associated with autonomous driving, lane-changing decisions, which involves a core model that determines whether a vehicle should change lanes based on its state, environment and traffic rules, are prominent. This model also determines the best strategy, aiming to ensure safe and efficient manoeuvres that adapt to current traffic conditions and improve overall performance[6].
@@
Figure 1: The importance of intent prediction and cooperative driving. Figure 1(a) highlights the importance of intent prediction. As autonomous vehicles approach an intersection, the ability to accurately predict the intentions of all nearby vehicles becomes critical. This capability enables the autonomous vehicle to make optimal decisions, ensuring both safe passage through the intersection and efficient manoeuvring while yielding to surrounding vehicles. Without intent prediction, autonomous vehicles would lose an important basis for decision making. This absence can affect the efficiency of crossing intersections and even lead to accidents. Figure 1(b) highlights the importance of cooperative driving for autonomous vehicles. The scenario shows three autonomous vehicles navigating an intersection simultaneously, with potential conflicts between their intentions. A unified coordination mechanism between these agents allows them to obey traffic rules or resolve conflicts efficiently. For example, in the current situation, the vehicle travelling straight ahead would have the right of way, followed by the autonomous vehicle closest to its target lane, and finally the more distant autonomous vehicle. Without coordination, maintaining safe distances and speeds could prevent a smooth intersection crossing.
@@
To address the above issues, we propose the Mix Q-learning for Lane Changing (MQLC) model. First, we achieve more accurate predictions of neighbouring vehicles’ intentions through deep learning-based trajectory prediction, outperforming current techniques. In addition, the classical Deep Q-Network (DQN) structure is extended to adapt to the lane-changing decision environment, enabling superior information acquisition and processing at the agent level, thereby facilitating the learning of safe and efficient lane-changing strategies. At the decision level, we plan to adopt a collaborative approach between agents in the context of CVIS, using joint decision making with both global and individual coordination. Specifically, We adopt the approach proposed by Zhang et al. [30], known as the ”QCOMBO” framework, to establish individual and global Q-networks. This framework aims to align the collective Q-decisions of all individuals with the global Q-estimation, thereby incorporating the global context into individual decision-making processes. However, in this framework, individual Q networks were still solely responsible for decision making without taking into account collaborative considerations. To address this limitation and to account for the collaborative nature of lane-changing decisions, we analyse the characteristics of collaboration in lane-changing decisions and introduce the paradigm of individual Q-network action preselection followed by global Q-network arbitration. This paradigm ensures that all agents make coordinated lane-changing decisions while maximizing overall utility and preserving individual gains. By enhancing both decision information analysis and decision formulation aspects, our method aims to achieve rational individual control and improve overall traffic efficiency. By enhancing these aspects, our model aims to increase the credibility and safety of the decisions made.
@@
2. We have achieved intention prediction based on deep neural networks that can better understand the interactions between vehicles, thereby enriching the decision process with valuable information. In addition, we improve the structure of the reinforcement learning network by combining the advantages of multiple neural networks to make it more suitable for lane-changing decision tasks. This ensures the decision network’s ability to make more informed and efficient lane-change decisions.
@@ II-A Rule-based models
Rule-based models aim to use predefined sets of rules and heuristics to guide lane-changing decisions. Their creation and design is based on expert knowledge. Rule-based models have attracted considerable attention from researchers due to their ease of application and the high interpretability of their decision processes [31]. Within this category, classic models include the City Driving Decision Structure introduced by Gipps et al.[7] and the generalised vehicle-following model for urban traffic further improved by Kesting et al.[8]. Other modelling methods include the use of Traffic Cellular Automaton (TCA) models, as demonstrated by Maerivoet et al.[11], and driver behaviour classification methods, as explored by Sun et al.[10]. These modelling methods help to better simulate lane changing behaviour in urban traffic environments.
However, rule-based models rely on expert knowledge and suffer from limited flexibility and generalisation, making them unsuitable for complex and dynamic road scenarios. As a result, researchers have shifted their focus to learning-based models.
@@
Supervised learning-based models typically use deep neural networks to regress human vehicle trajectory data to learn lane-changing decisions similar to human drivers. In recent years, supervised learning methods have emerged to address lane-changing problems, such as the use of Long Short-Term Memory (LSTM) neural networks combined with Conditional Random Fields (CRF) to learn human-like manoeuvring decisions[14], and the use of Bayesian parameter optimisation and Support Vector Machine (SVM) techniques[15]. To enrich limited decision information, Zhang et al.[13] integrated human factors represented by driving style into the study of autonomous lane changing (DLC) to achieve more rational lane changing decisions.
While supervised learning methods have achieved some success, they face significant challenges related to the collection and processing of human driver trajectory data, hindering the development of such methods. In addition, human lane-changing decisions are based on limited observations, and personality traits of human drivers can negatively influence these decisions. As a result, researchers have shifted their focus to reinforcement learning-based lane-changing decision methods.
@@
In recent years, several methods have been proposed to improve the performance of Deep Reinforcement Learning (DRL)-based lane-changing models. Based on the innovations and contributions of these methods, existing DRL-based lane-changing models can generally be classified into three types.
The first category focuses on the input of the model, aiming to improve performance by designing the observation space or incorporating specific network structures at the input end. For example, introducing an attention mechanism at the input end[32, 21] or adding a Bird’s Eye View (BEV) image with vector input[21] to improve the model’s information extraction capabilities. More recently, there has been a method that uses observed adversarial research to improve the robustness of the lane-changing decision model[33].
The second category emphasises the reward function, using cleverly designed reward functions to improve performance. Recently, there have been innovative ways to introduce a global traffic flow[34] or risk term[17] into the reward function to make the lane-changing model more harmonious and safer when changing lanes. Yuan et al.[35] decomposed the traditional single reward function into three sub-rewards, simplifying the learning difficulty of the lane-changing task.
The third category focuses on the algorithm and framework of the lane-changing model. Xu et al.[36] proposed a multi-objective approximate policy iteration algorithm to improve the safety, efficiency, and smoothness of lane-changing behaviour for autonomous vehicles. To add an extra layer of safety, the researchers incorporate structures such as safety verification[37] or safety supervisors[38] into the algorithmic framework to ensure that the model runs smoothly in real-world situations.
MQLC is a reinforcement learning-based lane-changing decision method that simultaneously improves both input and framework aspects. Compared to other reinforcement learning methods, MQLC achieves explicit coordination between lane-changing agents through additional inputs and a collaborative decision-making process.
@@ III-A Problem formulation
Based on the previous statement, we address the lane-changing decision problem in a multi-agent scenario. Our goal is to improve information acquisition, decision networks, and decision architecture to enable agents to achieve better operating conditions through more rational lane-changing strategies.
@@
In this paper, we consider two types of decision entities. While the global highway system has access to the complete state, the individual vehicles within the system only have access to partial information from $s_{t}$. Therefore, we formulate the multi-agent lane-changing decision problem as a Partially Observable Markov Decision Process (POMDP). In addition to the components of a regular MDP, such as the state space $\mathcal{S}$, the action space $\mathcal{A}$, the reward function $\mathcal{R}$, and the state transition model $\mathcal{P}$, POMDP introduces an observation space $\Omega$ ($o_{t}\in\Omega$) and an observation probability distribution $\mathcal{O}$,$\left(o_{t}\sim\mathcal{O}\left(s_{t}\right)\right)$ representing the observations in state $s_{t}$. Based on the POMDP model, each vehicle independently samples observations from the state and autonomously executes actions, receiving rewards provided by the environment, to progressively discover the strategy that can achieve the highest future rewards.
In this concept, the state transition model $\mathcal{P}$ is defined by the environment. Therefore, in the following a detailed introduction to the observation space $\Omega$, the state space $\mathcal{S}$, the action space $\mathcal{A}$ and the reward function $\mathcal{R}$ is given, which are essential for the decision making of the agents.
III-A1 Observation space
@@
The observation space is the collection of observations on which the vehicle’s decision making is based. Its design must accurately reflect the vehicle’s motivation for lane changing and the conditions in the environment that affect lane changing decisions. As mentioned earlier, the observation at time t for our vehicle consists of two parts: kinematic observations ($\Omega_{K}$) and driving intentions ($\Omega_{I}$).
(a) Kinematic observation: $\Omega_{K}$ are designed to represent the positional information and traffic conditions of the surrounding vehicles, allowing the ego vehicle to understand the environmental information and make lane-changing decisions that can improve its own traffic conditions. $\Omega_{K}$ serve as the basic observations on which the agent relies to make lane-changing decisions.
For agent $i$ at time $t$ with perception distance $P_{D}$, its kinematic observation can be represented as a matrix $K_{i}^{t}$ of size ${V}\times{F}$, where $V$ is the number of vehicles observed and $F$ is the number of attributes for each vehicle. In this paper we consider five attributes that the agent can obtain through kinematic observations, including attributes to distinguish agents at offset 0 from non-existent agents, the world offset of the ego vehicle on the x-axis and the y-axis, the velocity on the x-axis and the y-axis of the vehicle, denoted as presence, x, y, $v_{x}$ and $v_{y}$ respectively.
(b) driving intentions: $\Omega_{I}$ aims to enable the ego vehicle to consider the potential future driving actions of surrounding vehicles when making decisions.
Zhang et al.[22] use predicted trajectories to represent driving intentions, which has clear advantages in terms of lane-changing behaviour and right-of-way considerations. While their approach captures essential characteristics of driving intent, it has limitations. They use a simplified polynomial-based trajectory prediction method for surrounding vehicles, treating them as driving intent added to observations for decision making. But the methods often fail to accurately reflect the true intent of surrounding vehicles, leading to errors that can render intent predictions ineffective for decision making.
@@
Based on the above analysis, we propose a novel approach to intent prediction. We encode the scene from the previous time step $T_{P}$ and feed it into a neural network to predict the coordinates of the surrounding vehicles at a future time step $T_{F}$. This approach allows a more accurate estimation of intentions by anticipating future vehicle positions. Compared to polynomial-based methods, deep learning-based trajectory prediction offers stronger capabilities in non-linear modelling, adaptability and generalisation. It can extract more complex trajectory patterns and dynamic behaviours from trajectory data and adapt to changing driving environment conditions, thereby maintaining higher prediction accuracy. Considering the requirements of the lane-changing decision model in terms of parameter set and computational efficiency, this paper plans to use a lightweight trajectory prediction network, and the specific network architecture is shown in Table I of the experimental section. The reason for choosing this architecture is that the vehicle coordinate trajectory is input into the network in the form of points. GCN can better extract the topological structure features of the vehicle, while GRU is used for time series trajectory prediction.
III-A2 State space
Since the state space serves as a dependent condition for global agents to make joint action decisions, its concept also needs to be clarified here.
The setting of the state space in this paper must satisfy two basic conditions: first, the state space should contain essential information for making joint action decisions; second, the state space should reflect as much key information from the environment as possible. Based on the above analysis, we adopt the set of observation spaces of all agents as the state space. The shape of the state space can be represented as a matrix of size $T_{n}$$\times$×F, where $T_{n}$ represents the number of vehicles that the entire state space can contain, which can be expressed by $n$$\times$ $V$. $V$ and $F$ have the same meaning as in the previous text, and $n$ represents the number of agents. This design is based on several considerations. First, the state space is intended for joint action estimation of the global agents. By using the concept of an observation space set, it ensures a comprehensive understanding of the environment, which allows a reasonable benefit estimation for joint actions. Second, the joint actions are derived from the optimal actions sampled by each vehicle agent based on its own observations which promotes effective learning in subsequent stages. Finally, this state space configuration essentially concatenates the observation spaces, satisfying the $o_{t}\sim s_{t}$ requirements in its form.
III-A3 Action space
@@
For vehicle agent $i$ at time $t$, the action $a_{t}^{i}$ it takes is sampled from its advantage estimation over all feasible actions in the discrete action space. The lane-changing behaviour discussed in this paper refers to all actions involved in controlling the vehicle for normal driving. Therefore, the action space $\mathcal{A}$ consists of five discrete actions:
$a_{t}\in\mathcal{A}=\{lane_{left},idle,lane_{right},faster,slower\},$ (1)
@@
Based on the definitions above, the joint action space is the set of all possible actions for all vehicles, represented in matrix form as $A_{joint}=A_{1}\times A_{2}\times\cdots\times A_{n}$, where $n$ is the number of agent vehicles. The joint action space $A_{joint}$ contains the complete set of actions that all vehicles in the environment can possibly take at a given time step.
III-A4 Reward function
The reward refers to the signal given to the agent after it has performed an action and the environment changes based on the state transition model. Rewards play a crucial role in guiding the agent’s training strategy in reinforcement learning.
In the lane-changing decision task, the design of rewards must take into account the rationality, safety, and efficiency of the lane-changing behaviour. Therefore, the design of the reward function in this paper is based on the following objectives: First, safety is one of the most critical criteria for evaluating lane-changing behaviour. Second, lane-changing behaviour often involves compromise and cooperation. Unnecessary lane changes can create avoidable risks for all traffic. Finally, lane-changing behaviour is motivated by the desire to improve the vehicle’s operating conditions, such as better visibility or more buffer space. In particular, a reasonable lane change should result in an increase in the vehicle’s speed. Therefore, speed improvement is another aspect that needs to be considered in the reward design. This paper hopes to verify the effectiveness of the proposed collaborative method in lane-changing decision-making, so the default reward items and functions in the virtual environment are selected.
Based on the above analysis, the proposed reward function $\mathcal{R}$ takes the following form:
@@ III-A4 Reward function
where $R_{1}$, $R_{2}$, and $R_{3}$ represent the safety reward term, unnecessary lane change penalty term, and high-speed reward term mentioned earlier, respectively. $\omega_{1}$, $\omega_{2}$, and $\omega_{3}$ are the respective coefficients for these three reward terms. Below is the explanation of the specific settings for each reward term. The safety reward term $R_{1}$ aims to penalize lane-changing behaviors that jeopardize driving safety. Its formula is as follows:
$R_{1}=\left\{\begin{array}[]{ll}-1,&{Vehicle\quad crash}\\ 0,&{Otherwise.}\end{array}\right.$ (3)
The Unnecessary Lane-Changing Penalty term $R_{2}$ aims to discourage unnecessary lane-changing behaviour. Due to the limited observation range of vehicles and the uncertainty of the impact of current decisions on the future, it is difficult to determine the ’unnecessary’ nature of lane-changing actions. Instead, the $R_{2}$ term is designed to encourage all agent vehicles to stay in the rightmost lane. Its formula is as follows
$R_{2}=\frac{L_{c}}{L_{r}},$ (4)
where $L_{c}$ is the current lane number of the vehicle and $L_{r}$ is the lane number of the rightmost lane. Lane numbers are assigned based on the arrangement of lanes from left to right in the environment, with lane 1 being the leftmost lane and increasing incrementally to the right. This reward condition provides a higher reward for agents who are closer to the rightmost lane, encouraging them to prefer to drive in the rightmost lane and thus avoid unnecessary lane changes.
As for the fast reward term $R_{3}$, its formula is represented as follows
$R_{3}=\frac{v-v_{\min}}{v_{\max}-v_{\min}},$ (5)
where $v$, $v_{\min}$ and $v_{\min}$ represent the agent’s current speed, the set minimum speed and the set maximum speed respectively. The high speed reward encourages the vehicle to approach the maximum speed to ensure that the speed requirement for the reward is met.

### v2

@@ Abstract
Lane-changing decisions, which are crucial for autonomous vehicle path planning, face practical challenges due to rule-based constraints and limited data. Deep reinforcement learning has become a major research focus due to its advantages in data acquisition and interpretability. However, current models often overlook collaboration, which affects not only impacts overall traffic efficiency but also hinders the vehicle’s own normal driving in the long run. To address the aforementioned issue, this paper proposes a method named Mix Q-learning for Lane Changing (MQLC) that integrates a hybrid value Q network, taking into account both collective and individual benefits for the greater good. At the collective level, our method coordinates the individual Q and global Q networks by utilizing global information. This enables agents to effectively balance their individual interests with the collective benefit. At the individual level, we integrate a deep learning-based intent recognition module into our observation and enhance the decision network. These changes provide agents with richer decision information and more precise feature extraction for improved lane-changing decisions. This strategy enables the multi-agent system to learn and formulate optimal decision-making strategies effectively. Our MQLC model, through extensive experimental results, impressively outperforms other state-of-the-art multi-agent decision-making methods, effectively improving the overall efficiency. The code is available at https:github.com/pku-smart-city/source_code/tree/main/MQLC.
@@ I Introduction
In the field of CVIS, autonomous driving and related applications have the potential to revolutionise transport which could fundamentally change traffic patterns and provide significant convenience[2, 3, 4, 5]. Among the tasks associated with autonomous driving, lane-changing decisions, which involve a core model that determines whether a vehicle should change lanes based on its state, environment and traffic rules, are prominent. This model also determines the best strategy, aiming to ensure safe and efficient manoeuvres that adapt to current traffic conditions and improve overall performance[6].
@@
Figure 1: The importance of intent prediction and cooperative driving. Figure 1(a) highlights the importance of intent prediction. As autonomous vehicles approach an intersection, the ability to accurately predict the intentions of all nearby vehicles becomes critical. Figure 1(b) highlights the importance of cooperative driving for autonomous vehicles. This scenario shows that when there is a potential conflict between the intentions of the three autonomous vehicles, the unified coordination mechanism enables them to effectively obey traffic rules or resolve the conflict.
@@
To address the above issues, we propose the Mix Q-learning for Lane Changing (MQLC) model. First, we achieve more accurate predictions of neighbouring vehicles’ intentions through deep learning-based trajectory prediction, outperforming current techniques. In addition, the classical Deep Q-Network (DQN) structure is extended to adapt to the lane-changing decision environment, enabling superior information acquisition and processing at the agent level, thereby facilitating the learning of safe and efficient lane-changing strategies. At the decision level, we plan to adopt a collaborative approach between agents in the context of CVIS, using joint decision making with both global and individual coordination. Specifically, we adopt the approach proposed by Zhang et al. [30], known as the “QCOMBO” framework, to establish individual and global Q-networks. This framework aims to align the collective Q-decisions of all individuals with the global Q-estimation, thereby incorporating the global context into individual decision-making processes. However, in this framework, individual Q networks were still solely responsible for decision making without taking into account collaborative considerations. To address this limitation and to account for the collaborative nature of lane-changing decisions, we analyse the characteristics of collaboration in lane-changing decisions and introduce the paradigm of individual Q-network action preselection followed by global Q-network arbitration. This paradigm ensures that all agents make coordinated lane-changing decisions while maximizing overall utility and preserving individual gains. By enhancing both decision information analysis and decision formulation aspects, our method aims to achieve rational individual control and improve overall traffic efficiency.
@@
2. We have achieved intention prediction based on deep neural networks that can better understand the interactions between vehicles, thereby enriching the decision process with valuable information. In addition, we improve the structure of the reinforcement learning network by combining the advantages of multiple neural networks to make it more suitable for lane-changing decision tasks. This ensures that the decision network can make more intelligent lane change decisions and improve overall traffic efficiency.
@@ II-A Rule-based models
Rule-based models aim to use predefined sets of rules and heuristics to guide lane-changing decisions. Their creation and design are based on expert knowledge. Rule-based models have attracted considerable attention from researchers due to their ease of application and the high interpretability of their decision processes [31]. Within this category, classic models include the City Driving Decision Structure introduced by Gipps et al.[7] and the generalised vehicle-following model for urban traffic further improved by Kesting et al.[8].
However, rule-based models rely on expert knowledge and suffer from limited flexibility and generalisation, making them unsuitable for complex and dynamic road scenarios.
@@
Supervised learning-based models typically use deep neural networks to regress human vehicle trajectory data to learn lane-changing decisions similar to human drivers. In recent years, supervised learning methods have emerged to address lane-changing problems, such as the use of Long Short-Term Memory (LSTM) neural networks combined with Conditional Random Fields (CRF) to learn human-like manoeuvring decisions[14], and the use of Bayesian parameter optimisation and Support Vector Machine (SVM) techniques[15].
While supervised learning methods have achieved some success, they face significant challenges related to the collection and processing of human driver trajectory data. In addition, human lane-changing decisions are based on limited observations, and personality traits of human drivers can negatively influence these decisions. As a result, researchers have shifted their focus to reinforcement learning-based lane-changing decision methods.
@@
Based on the innovations and contributions of these methods, existing DRL-based lane-changing models can generally be classified into three types. The first category focuses on improving performance by enhancing the input of the model, such as using attention mechanisms[32, 21] or adding Bird’s Eye View (BEV) images with vector[21] input to boost information extraction capabilities. The second category improves performance by designing better reward functions. Recent approaches include incorporating global traffic flow[33] or risk terms[17] into the reward function to make lane-changing models more harmonious and safer. The third category emphasizes the algorithm and framework, adding safety layers like safety verification[34] or safety supervisors[35] to ensure smooth operation in real-world situations.
MQLC is deep a reinforcement learning-based lane-changing decision method that simultaneously improves both input and framework aspects. It achieves explicit coordination between lane-changing agents through additional inputs and a collaborative decision-making process.
@@ III-A Problem formulation
Based on the previous statement, we address the lane-changing decision problem in a multi-agent scenario. Our goal is to improve information acquisition, decision networks, and decision architecture to enable agents to achieve better operating conditions through more rational lane-changing strategies. We formulate the multi-agent lane-changing decision problem as a Partially Observable Markov Decision Process (POMDP) follow[22]. In addition to the components of a regular MDP, such as the state space $\mathcal{S}$, the action space $\mathcal{A}$, the reward function $\mathcal{R}$, and the state transition model $\mathcal{P}$, POMDP introduces an observation space $\Omega$ ($o_{t}\in\Omega$) and an observation probability distribution $\mathcal{O}\left(o_{t}\sim\mathcal{O}\left(s_{t}\right)\right)$ representing the observations in state $s_{t}$. Based on the POMDP model, each vehicle independently samples observations from the state and autonomously executes actions, receiving rewards provided by the environment, to progressively discover the strategy that can achieve the highest future rewards.
@@
Observation space: The observation space is the collection of observations on which the vehicle’s decision making is based. As mentioned earlier, the observation at time t for our vehicle consists of two parts: kinematic observations $\Omega_{k}$ and driving intentions $\Omega_{I}$.
Kinematic observation: $\Omega_{k}$ represents the kinematic state information of the ego vehicle and its surrounding vehicles, denoted as $K_{i}^{t}$ with a size of VV×$F$. Here, $V$ represents the number of observable vehicles, including the ego vehicle itself. $F$ corresponds to the set of vehicle features, represented as presence, $x$, $y$, $v_{x}$, and $v_{y}$, where the five values indicate whether the vehicle is alive, its position and velocity.
@@
Driving intentions: $\Omega_{I}$ aims to enable the ego vehicle to consider the potential future driving actions of surrounding vehicles when making decisions. We propose a novel approach for intention prediction. We encode the scene from the previous time step $T_{P}$ and feed it into a neural network to predict the coordinates of the surrounding vehicles at a future time step $T_{F}$ . Compared to polynomial-based methods[22], deep learning based trajectory prediction offers stronger capabilities in nonlinear modelling, adaptability and generalisation. The specific network architecture is shown in Table I of the experimental section. The reason for choosing this architecture is that the vehicle coordinate trajectory is input into the network in the form of points. GCN can better extract the topological structure features of the vehicle, while GRU is used for time series trajectory prediction.
@@
Action space: for vehicle agent i at time t, the action $a$ it takes is sampled from its advantage estimation over all feasible actions in the discrete action space. And the action space $\mathcal{A}$ consists of five discrete actions:
$a_{t}\in\mathcal{A}=\{lane_{left},idle,lane_{right},faster,slower\}.$ (1)
@@
The joint action space is the set of all possible actions for all vehicles, represented in matrix form as $A_{joint}=A_{1}\times A_{2}\times\cdots\times A_{n}$, where n is the number of agent vehicles.
Reward function[36]: The reward is a signal provided to the agent after it takes an action. This signal reflects changes in the environment resulting from the state transition model. This paper hopes to verify the effectiveness of the proposed collaborative method in lane-changing decision-making, so the default reward items and functions in the virtual environment[36] are selected.
@@
where $R_{1}$, $R_{2}$, and $R_{3}$ represent the safety reward term, unnecessary lane change penalty term, and high-speed reward, respectively. $\omega_{1}$, $\omega_{2}$, and $\omega_{3}$ are the respective coefficients for these three reward terms.